OpenStudy (anonymous):

10.8 is 60% of what number?

OpenStudy (anonymous):

60% of some number x is the same as saying 0.6 * x. So, you have 10.8 = 0.6 * x. Solve for x by multiplying each side by 1/ 0.6. Then you have 10.8 / 0.6 = x.